### v2

# Strain Induced Kramers-Weyl Phase in III-V Zinc Blende Systems

Abstract
We present theoretical observations on the topological nature of strained III-V semiconductors. By $k\cdot p$ perturbation, it can be shown that the strain-engineered conduction band hosts a Kramers-Weyl node at the $\Gamma$ point. It is theoretically shown a curated strain can create and then tune the sign of the topological charge. Furthermore, we outline experimental methods for both the realization and detection of strain-induced topological phase transitions.
preprint: APS/123-QED
I Introduction
Topological condensed matter systems highlight the intersection between fundamental physics and practical technologies. On the fundamental side we observe emergent quasi-particles such as Dirac or Weyl fermions; with them we can explore exotic physics such as topological charge transport and chiral anomaly[1]. The practicality manifests through motivations to increase the efficiency of modern electronics via improving resistance-product areas[2], low energy switching in spintronics [3], and the creation of impurity resistant devices. This further drives innovations in the design and engineering of topological quantum computing platforms such Majorana zero modes and fractal quantum hall states [4, 5].
Mirror symmetry can be removed by the application of curated strains. For example, in Fig. 1A, a pristine crystal–marked in dashed lines–has two mirror symmetry elements m’; by applying a uniaxial planar force along any directions that are not x,z, nor $<$101$>$, one can remove these two mirror symmetry elements. Such forces can be generalized to 3D as shown in Fig. 1B. In the example of cubic zinc blende (ZB), the crystal lacks inversion symmetry in its natural state, tuning the pictured $\theta$ and $\phi$ breaks the mirror symmetry making the crystal chiral.
Figure 1: A) The removal of symmetry elements m’ due to a distortion in the crystal. B) A 3-Dimensional force as the generator of the crystallographic distortion due to strain.
II Zinc Blende as a case study
A good case study is the cubic ZB Crystal system which, as opposed to diamond, lacks an inversion-center. The spin-orbit coupling in such bulk semiconductors can be written as a sum of the contributions of its inherent inversion-asymmetry[6] plus a strain induced component: [7, 8, 9].
$H=H_{0}+\Omega_{Dress}\cdot\sigma+\Omega_{Strain}\cdot\sigma$ (1)
$\Omega_{Dress}^{x}=\gamma\textit{k}_{x}(\textit{k}_{y}^{2}-\textit{k}_{z}^{2})$ (2)
$\Omega_{Strain}^{x}=C(\epsilon_{xy}\textit{k}_{y}-\epsilon_{xz}\textit{k}_{z})+D\textit{k}_{x}(\epsilon_{zz}-\epsilon_{yy})$ (3)
where $H_{0}=(\hbar\textit{k})^{2}/2m$, and $\epsilon_{ij}$ is the symmetric strain tensor.
In the above, $\gamma$, C, and D are material constants[10]. The prefactor $\gamma$ ,commonly called the Dresselhuas coefficient, reflects the cubic in k spin-orbit coupling strength of ZB crystals. The influence of the second term $\Omega_{Strain}$ has been studied extensively by Seiler et al. via Shubnikov de Haas oscillations and showed both experimentally and numerically the influence of a few kilobar of pressure on the Fermi surface and conduction band-splitting in InSb[11]. Further discussion on the C, and D coefficients can be found in [10, 12, 8].
Figure 2: The influence of mirror symmetry-breaking strain on band structure and spin-texture. A) A diagram showing the displacement of atoms when acted on by an exaggerated force. B) A general ZB band structure with spin degeneracy along $\Gamma L$ and $\Gamma X$ when mirror symmetry is present, and spin-split when broken. C-D) The spin texture painted on a uniform sphere with an arbitrary radius in k-space with care taken to respect any approximations necessitating small k values. The spin texture on the k-sphere is show as arrows and, for emphasis, the spin texture perpendicular to the surface is drawn as the colormap. Of the two, the left sphere is presented as an unstrained comparison and reflects precisely Dresselhaus spin-orbit coupling behaviour. The rightmost sphere is the result of a force acting on a ZB crystal at $(\theta,\phi)=(\frac{\pi}{4},\frac{\pi}{6})$ when the coefficient C is suppressed.
In this letter we will examine the topological properties encoded within Eq 3. Before discussing the more exotic effects of strain, we point out that any force will have an effect on the severity of the bulk inversion-asymmetry, elevating the corresponding spin-orbit coupling strengths. Fig. 2 illustrates the action of strain on a ZB system by means of spin texture and band structure. In the top left, A) displays a cartoon of an exaggerated strain on the crystal structure. In B), the presence of mirror symmetry in the time reversal invariant ZB lattice is reflected by the spin degeneracy along $\Gamma-L$ and $\Gamma-X$ as seen in the band structure plots (top); the breaking of mirror symmetry results in the spin splitting when tracing the plotted k-vectors (bottom). The spheres in Fig. 2 are a visualization the spin texture of the same two systems with the colormap showing the normalized spin texture perpendicular to the surface. In the unstrained prototype, all spin vectors are parallel to the surface which remains true for a sphere of any radius centered at k= 0. Consequently, Dresselhaus spin orbit coupling does not contribute to the Chern number which is the Berry curvature flux through the Brillouin zone. Lastly, the strained spin texture is painted on the final sphere with the principal spin vectors presented as an inset. The flux is now not necessarily zero which makes non-zero topological charge a possibility. It is observed that the normal strains along the three principal axes are the key for leading to the inward or outwards spin textures. Inspecting the inset reveals that our bounding sphere encloses a positive topological charge. We can engineer $\mathcal{C}_{Weyl}$=-1,+1 as well as trivial states by manipulating the strain.
Reference [13] showed that any non-magnetric chiral crystal harbors a Kramers-Weyl node, coupled with knowledge that any crystal lacking an inversion center can always be strain engineered to be chiral[14], we have a path forward to the realization of topological states by straining otherwise trivial semiconductors.
The term in the Hamiltonian that describes a Weyl node is
$H_{Weyl}=A_{ij}k_{i}\sigma^{j}$ (4)
which has similar structure as Eqs (1,3), examining like terms, we can populate coefficients of a strain induced Weyl Hamiltonian. Explicitly,
$A_{ij}^{strain}=\begin{pmatrix}D(\epsilon_{zz}-\epsilon_{yy})&-C\epsilon_{yx}&C\epsilon_{zx}\\ C\epsilon_{xy}&D(\epsilon_{xx}-\epsilon_{zz})&-C\epsilon_{zy}\\ -C\epsilon_{xz}&C\epsilon_{yz}&D(\epsilon_{yy}-\epsilon_{xx})\end{pmatrix}$ (5)
Notice that the Kramers-Weyl node is completely characterized by elements of the the strain tensor. Furthermore, the topological charge can be evaluated given just Eq 5 via $\mathcal{C}_{Weyl}=SGN(DET(A))$[1, 13].
Uniaxial force shown in figure 1A will not remove all mirror symmetries of the lattice such as those on 001,010,100 and equivalent directions. We also expect uniaxial force along $<$110$>$and $<$111$>$ will also preserve the mirror symmetry of the crystal. We can thus expect the determinant of the Eq. 5 to vanish in such cases. Beyond these special test cases, it is not obvious which force vectors, $\mathbf{F(\theta,\phi)}$, transform the material into which topological state. As a means to explore this phase space, we create a topological phase diagram.
By considering a single material system, the 6 degrees of freedom in the strain tensor are reduced to 2, the azimuthal and polar angles $0\leq\phi\leq 2\pi$; and $\leq\theta\leq\pi$. A set compliance tensor will unambiguously transform some normalized force into a symmetric strain tensor. The direction of this force, described in spherical coordinates, alone defines the topological state of an otherwise topologically trivial crystal–as seen in Fig. 1. By normalized force, we mean a reasonable pressure maybe on the order of $10^{5}$kPa to $10^{6}$kPa being acted upon the crystal, flipping the direction of the force, flips the sign of the strain. The materials we investigate here include GaAs and InSb due to their relative popularity and moderate Young’s moduli making the experimental realization of larger strains more feasible. But the results apply for any III-V ZB crystals. We evaluate SGN(DET(A)) for the range of angles mentioned and display the resulting topological phase diagram in figure 3. Here we show the two main limiting cases of $C>D$ and $D>C$ in the top and bottom panel respectively. In physical situations there is dominance of the shear components which is owed to C being generally larger than the others. In the case of GaAs, C is an order of magnitude larger than D and D is on similar footing as $\gamma$ at $35eVA^{3}$[15]
Figure 3: Topological phase diagram of a ZB lattice being acted on by a force $\mathbf{F(\theta,\phi)}$. Top row is the limiting case of C $>$ D and the bottom row is for larger D. Small values are thresholded for numerical stability.
The nature of Fig. 3 can be interpreted by examining a few special cases. As previously mentioned, forces lying within the 110 plane and principal axes do not create a chiral structure; meaning that the models remain topologically trivial for forces along $\phi=(2n+1)\frac{\pi}{4}$ for all $\theta$, here n is an integer. Additionally, the Chern number vanishes for forces enacted along $sin(\phi)=\pm cot(\theta),cos(\phi)=\pm cot(\theta)$. These states correspond to forces which preserve the mirror symmetries by means of symmetric projections.
Strain engineering is a mature and developed research topic; in particular, epitaxial strain is a well-known and growing field for the strain engineering of materials[16, 17]. Today, around 90% of produced silicon transistor is strained to meet desired specifications[18]. Piezoelectric is another option if tunability is desirable, samples can either be mounted directly on top of piezoelectrics or be prepared in specialized devices.[19]. Regardless of the preferred method, there is also the issue of the maximal strain a bulk sample can sustain. Nanomembranes show a theoretical limit of around 10% [20] though bulk samples will be lower due to the increased likely hood of defects. Whether or not a maximal strain is sufficient depends on the material properties.
A difficulty adjacent to growing/creating these test systems is the characterization of the desired topological properties. In the case of most III-V semiconductors there are some liberties that can be taken; since we have a 1:1:1 tri-correspondence between the existence of the Kramers-Weyl node, mirror symmetry breaking, and chirality, the identification of any of these is equivalent to the detection of the others.
In the realm of the optical methods there exists Angle Resolved Photoemission Spectroscopy[21, 22] for the detection of topological states and Circular Photo-Galvanic effect measurements for the detection of chirality and Weyl system which has been demonstrated with the direct detection of Weyl nodes[23, 24]. Other possible avenues which have had substantial attention recently include Faraday rotation and Magneto Optical Kerr Effect experiments, chiral anomaly, and bilinear magnetoresistance measurements. All of these have already had bountiful discussion on literature so we will not further talk about them here[10, 25, 26, 27, 28]. Instead we would like to explore the strain-tuned quantum oscillations in III-V semiconductors[11]. Shubnikov de Haas oscillations are a direct measurement of the cross-sectional area of the Fermi surface[29, 30, 31]. If we take the magnetic field direction as the z-axis, the oscillation frequency is proportional to the square of the azimuthually averaged Fermi-vector. The warped-sphere model, characteristic of III-V systems, makes the tracking and fitting of the oscillations fairly straight forward. In this manner, if the angle between the surface normal of the sample and the magnetic field is scanned through a mirror symmetry plane one can deduce whether mirror symmetry is present or not. For example, in Fig. 4 we present a Monte Carlo simulation of a quantum oscillations experiment. A) is a general overview of device structure with the important parts being the 4-probe wiring and scanning angle off the sample normal, [001] in this case. InSb, a very common semiconductor, is frequently grown with (100) as a major facet and will be considered for this section. In this experimental configuration we can scan through the (110) mirror plane as well as symmetric projections on the plane and record the oscillation frequencies as a function of angle. The results for two programmatically generated Fermi-surfaces is then presented as C) for a full $2\pi$ rotation. The unstrained control sample is reasonably symmetric about $\bar{1}10$,and $01\bar{1}$ 0 just as in reference [11]. The sensitivity of the Fermi surface is apparent both in the band structure of Fig. 2 and the inset of Fig. 4. More notably, the oscillation frequency is not symmetric about crossings of mirror symmetry planes.
Figure 4: An experimental perspective. A)Diagram illustrating a 4-terminal angle resolved magneto-resistance measurement. The scanning angle is defined from the surface normal of the sample [001] and the magnetic field vector. B)First Brillouin Zone for Zinc Blende crystal structure with 4 symmetry points highlighted. C) A Monte Carlo simulation of quantum oscillation frequency vs angle for $\sigma=0$ (green) and $\sigma\neq 0$ with strain on the order of 0.1%(purple).
III Conclusion
We presented and discussed the prospects of strain engineering crystals lacking an inversion center into topological hosts of Kramers-Weyl node. The spin-texture, band structure and topological phase space for strained Zinc Blende is investigated and explained. Lastly, we presented the effect of the strain on the quantum oscillations for observing the desired effects. While we kept the scope of the conversation limited to materials already lacking an inversion center, we are optimistic this work can be extended to more general systems by strain engineering both inversion and mirror asymmetries.
IV Acknowledgment
We acknowledge the funding support from US AFOSR under award number FA9550-23-1-0310.